Session: Nov 1999 (session 4)
Wavelength: 6 cm


To give feedback on this experiment, click here

Experiment Feedback

Westerbork (Tony Foley at Tue Nov 23 14:35:07 MET 1999) Minor failures
Stopped observing early as all scans after 11:25 are out of Hour Angle range. Total power in array lost from 09:32 to 09:40 UT (delay sprang to test setting). Tape lost vacuum at 10:04 so the end of pass 6, all passes 7 and 8 are lost. Restarted with pass 9 at ~10:55UT. Cause of vacuum loss was dirt from the tape blocking vacuum sensor.

Medicina (mauro at Tue Nov 23 20:47:23 MET 1999) Success

Medicina (mauro at Tue Nov 23 20:51:39 MET 1999) Success
All scan recorded Sky clear

Effelsberg (A. Kraus at Wed Nov 24 12:02:16 MET 1999) Success

Medicina (C.Trigilio at Thu Nov 25 11:00:33 MET 1999) Success

Noto (C.Trigilio at Thu Nov 25 11:00:51 MET 1999) Success

Jodrell (Alastair Gunn at Wed Dec 1 03:38:15 MET 1999) Minor failures
Cryogenics began to fail and Tsys increases from about 16:30 UT. No other known problems.

Onsala (John Conway at Mon Dec 13 09:53:43 MET 1999) Minor failures
Apparently OK. $$$ on readback tests though.

Calibration (Jean\-Francois Desmurs at Wed Dec 15 13:16:41 MET 1999) Success
The ANTAB formatted file for this experiment can be found on VLBEER in the nov99 directory under the name: gb035b.antab. Read the comments in the file to get further information.